Jeb has 15 coins. They are all quarters and dimes. He has a total of $2.10. How many dimes does he have?
Answer by Alan3354(21605)

Jeb has 15 coins. They are all quarters and dimes. He has a total of $2.10. How many dimes does he have?
-----------------
D = # of dimes
Q = # of quarters
D+Q = 15
10D + 25Q = 210
Q = 15-D (from the 1st eqn)
Sub into the 2nd eqn
10D + 25*(15-D) = 210
10D + 375 - 25D = 210
-15D = -165
D = 11
To check:
Q = 15-D = 4
4*25 + 11*10 is 210, so it's correct.